$SPAD/src/input e02bdf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
ncap7:=14
lamda:Matrix SF:=
 [[0.0 ,0.00 ,0.00 ,0.00 ,1.00 ,3.00 ,_
   3.00 ,3.00 ,4.00 ,4.00 ,6.00 ,6.00 ,6.00 ,6.00 ]]
c:Matrix SF:=
 [[10.00 ,12.00 ,13.00 ,15.00 ,22.00 ,26.00 ,_
   24.00 ,18.00 ,14.00 ,12.00 ,0.00 ,0.00 ,0.00 ,0.00 ]]
result:=e02bdf(ncap7,lamda,c, -1)

———-

References^{[1] nothing}

4